# Duistarmaat’s triple index and the difference of Hermitian matrices

### v1

@@ 1. Introduction
An important problem in the spectral theory of metric graphs is to quantify changes in the Laplacian eigenvalues when a finite number of vertex conditions are altered. [^1] For a survey of results when the vertex conditions are of a particularly well-studied type, we refer the reader to [5, Sec. 3.1]. Abstractly, this amounts to comparing self-adjoint extensions of a symmetric operator with finite and equal defect numbers. It was recently observed [4] that the difference of eigenvalue counting functions can be efficiently bounded in terms of a certain topological index.
The relevant index, introduced by Duistermaat in [9], takes three Lagrangian subspaces as its input. In the metric graph setting these correspond to the two sets of vertex conditions under consideration and also the Dirichlet conditions at the affected vertices. Duistarmaat’s definition is somewhat indirect, as it involves an arbitrary choice of a fourth Lagrangian plane that is transversal to the three planes of interest. The proof that the resulting index is independent of this choice relies on the Maslov index for paths of Lagrangian subspaces, see [9, Eq. (2.16)]. The Maslov index is an invariant depending on the homotopy class of the path (not just its endpoints), and hence it is inherently more complicated than an index depending on fixed Lagrangian subspaces.
@@
The above index takes values between 00 and $n$:
@@
To motivate the terminology in (3), we recall the coboundary operator $\delta$ (see [11, Chapter 3]) which maps functions of $k-1$ arguments to functions of $k$ argument by
@@
The plan for the rest of the note is as follows: in Section 2 we show that an index satisfying the conditions in Theorem 1.1 exists and in Section 3 we establish some further properties of this index. In Section 4 we interpret the Duistermaat index as the minimal value of the Maslov index [^2] Our definition of the Maslov index, which is reviewed in Section 4, follows the normalization convention of [20]; see [4, Sec. A.4] for a comparison with other conventions in the literature. over non-decreasing paths, namely
@@
Finally, in Section 5, we prove Theorem 1.4, which we use to deduce (1.1) and its extensions from equation (1.2). We also give an elementary derivation of (1.1) using the Haynsworth inertia formula.
@@ Acknowledgements
This material is based upon work (G.B.) supported by a grant from the Institute for Advanced Study (IAS), School of Mathematics. G.B. also acknowledges support from the NSF Grant DMS-2247473 and from the Association of Former Students of Texas A&M University (through the Faculty Development Leave program that supported the leave to IAS). G.C. acknowledges the support of NSERC grant RGPIN-2017-04259. Y.L. was supported by the NSF grant DMS-2108983/2106157, and would like to thank the Courant Institute of Mathematical Sciences at NYU and especially Prof. Lai-Sang Young for their hospitality. S.S. was supported in part by the NSF Grant DMS-2418900, by the Simons Foundation grant MP-TSM-00002897, and by the Office of the Vice President for Research & Economic Development (OVPRED) at Auburn University through the Research Support Program Grant.
@@ 2. Existence of the Duistermaat index
$E(\mathcal{M}):=\{\epsilon\in\mathbb{R}:X+\epsilon Y\text{ is not invertible}\}$ (2.4)
@@
$\𝐸R^{\epsilon}:=Y(X+\epsilon Y)^{-1},\qquad\epsilon\in\mathbb{R}\backslash E(\mathcal{M}),$ (2.5)
@@
is Hermitian, therefore $R^{\epsilon}$ is Hermitian for all $\𝐸\epsilon\in\mathbb{R}\backslash E(\mathcal{M})$.
@@
We follow the proof of a similar identity in [16, Lem. 8.3] (attributed therein to Brian Feldstein). We first compute the form $\Omega$ defined in (2.1) as a Hermitian form on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}\oplus\mathbb{C}^{n}$, where each $\mathbb{C}^{n}$ parametrizes the corresponding Lagrangian plane $\mathcal{L}_{j}$ via a frame $(X_{j},Y_{j})$. If, for instance,
@@
=$=\big{\{}v\in\mathbb{C}^{n}\colon 0=\left<u,(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k})v\right>_{\mathbb{C}^{n}}\text{ for all }u\in\mathbb{C}^{n}\big{\}}$
@@
Finally, to estimate the range of the possible values of the index, we use a symplectic transformation $\mathsf{S}$ such that $\mathsf{S}\mathcal{L}_{3}=\mathcal{G}_{\infty}$ to reduce the question to bounding the range of $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{G}_{\infty})$. That the latter is between 00 and $n$ follows from a version of equation (2.7):
@@
Equation (2.21) was already established in [4, Eq. (3.32)], but we repeat the proof here for completeness. We use equation (2.7) with $R^{\epsilon}_{3}=\epsilon^{-1}I$, and the desired result follows once we establish that $n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)=0=n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)$. In other words, we claim that for any Lagrangian frame $(X,Y)^{T}$, the Hermitian operator $\epsilon^{-1}I-Y(X+\epsilon Y)^{-1}$ is non-negative definite for small enough $\epsilon>0$. It is equivalent to consider
@@ 3.1. General identities
00 0 $=(\delta\!\operatorname{\iota})(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ (3.9)
@@ 3.2. Self-adjoint linear relations
For a given splitting of the symplectic space (in our case, $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$), every Lagrangian plane can be viewed as a self-adjoint linear relation; see [18, Sec. 14.1] or [7, Sec. 4.2]. In this context, the difference of two Lagrangian planes, $\mathcal{L}$ and $\mathcal{M}$, is the Lagrangian plane
:=$\mathcal{L}-\mathcal{M}:=\left\{\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}-v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}\end{pmatrix}\in\mathcal{L},\ \begin{pmatrix}u\\ v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathcal{M}\right\},$ (3.11)
@@
$\operatorname{dom}\mathcal{L}$ :=$:=\left\{u\in\mathbb{C}^{n}:\begin{pmatrix}u\\ v\end{pmatrix}\in\mathcal{L}\text{ for some }v\in\mathbb{C}^{n}\right\},$ (3.13)
$\operatorname{mul}\mathcal{L}$ :=$:=\left\{v\in\mathbb{C}^{n}:\begin{pmatrix}0\\ v\end{pmatrix}\in\mathcal{L}\right\},$ (3.14)
@@
=$\mathcal{L}=\big{\{}(x,Lx+y)\colon x\in\operatorname{dom}\mathcal{L},\ y\in\operatorname{mul}\mathcal{L}\big{\}}$ (3.15)
for some Hermitian matrix $L\colon\operatorname{dom}\mathcal{L}\to\operatorname{dom}\mathcal{L}$, see, e.g. [2, Thm. 1.5.1] or [18, Prop. 14.2].
@@
$\mathcal{L}-\mathcal{G}_{A}$ =$=\big{\{}(x,Lx-Ax+y)\colon x\in\operatorname{dom}\mathcal{L},\ y\in\operatorname{mul}\mathcal{L}\big{\}}$ (3.22)
=$=\big{\{}(x,(L-PA)x+y^{\prime})\colon x\in\operatorname{dom}\mathcal{L},\ y^{\prime}\in\operatorname{mul}\mathcal{L}\big{\}}.$
@@ 4.1. The Kashiwara index
Another commonly used index of a triple of Lagrangian planes arose independently in the work of Hörmander, Kashiwara and Wall; see, for instance, [19], [13, Eq. (3.3.15)], [14, Def. 1.5.1], [8, Sec. 8], [10, Sec 4.2]. The Hörmander–Kashiwara–Wall index $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ is defined [^4] Only the latter reference considered complex symplectic space; the signs in our definition of the form $\Omega$ are different. as the signature of the form $\Omega$ from Theorem 2.1,
@@
Proposition 4.2 is well-known when (4.7) is replaced by the condition $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=s(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$, see, e.g. [16, Thm. 8.4(d)] or [1, Prop. 4.4]. But the indices $s$ and $\operatorname{\iota}$ are related by an invertible linear transformation involving the invariants appearing in (4.5). ∎
@@
While usually established before the more complicated case of a triple, see [16, Exercise 8.1], this fact also follows from Proposition 4.2 by applying it to the triples $(\mathcal{L}_{1},\mathcal{L}_{1},\mathcal{L}_{2})$ and $(\mathcal{M}_{1},\mathcal{M}_{1},\mathcal{M}_{2})$ and using (3.1). ∎
@@ 4.2. The Maslov index
Suppose $\mathcal{L}\colon[0,1]\to\Lambda(n)$ is a differentiable path of Lagrangian planes, given by a frame $\big{(}X(t),Y(t)\big{)}$. For $t_{0}\in[0,1]$ and $v\in\mathcal{L}(t_{0})$ we define the [^5] The equivalence of (4.9) to other common formulas in the literature is proved in [4, Thm. 2.1]. crossing form $\mathfrak{q}$ on $\mathcal{L}(t_{0})$ by
@@
To calculate this Maslov index we define the two-parameter family $\mathcal{L}(s,t)=\mathcal{G}_{tQ-sI}$. It is clear that $\mathcal{L}(t)=\mathcal{L}(0,t)$ is homotopic, with fixed endpoints, to the concatenation of $\mathcal{L}(s,0)|_{s\in[0,s_{*}]}$, $\mathcal{L}(s_{*},t)|_{t\in[0,1]}$ and $\mathcal{L}(s_{*}-s,1)|_{s\in[0,s_{*}]}$, so the index of $\mathcal{L}(t)$ is the sum of these three indices. We choose the parameter $s_{*}>0$ to be small enough to guarantee that neither $M$ nor $M-Q_{\operatorname{dom}\mathcal{M}}$ has any spectrum in the interval $[-s_{*},0)$.
@@
The path $\mathcal{L}(s,0)|_{[0,s_{*}]}$ may still have a conjugate point at $s=0$. However, the $s$-crossing form, computed via (4.9), is $x,x\right>_{\mathbb{C}^{n}}$, which is negative definite and hence generates no contribution to (4.10) at the initial point of a path. We conclude that $\operatorname{Mas}_{[0,s_{*}]}\!\big{(}\mathcal{L}(s,0),\mathcal{M}\big{)}=0$ and, by similar reasoning, $\operatorname{Mas}_{[0,s_{*}]}\!\big{(}\mathcal{L}(s_{*}-s,1),\mathcal{M}\big{)}=0$.
We now evaluate the remaining term, $\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(s_{*},t),\mathcal{M}\big{)}$. The $t$-crossing form evaluates to $x,Qx\right>_{\mathbb{C}^{n}}$, which we claim to be strictly positive on $\mathcal{L}(s_{*},t)\cap\mathcal{M}$ whenever this intersection is non-trivial. Indeed, the form is obviously non-negative and if $Qx=0$ for some nonzero $x\in\operatorname{dom}\mathcal{L}$, then (4.16) implies $Mx=-s_{*}x$, which is excluded by our condition that $M$ has no spectrum in $[-s_{*},0)$.
@@
As a corollary of Lemma 4.6 we obtain the Zhou–Wu–Zhu identity [20, Thm. 1.1].
@@
Use Lemma 4.6 to choose a path ~$\widetilde{\mathcal{L}}\colon[1,2]\to\Lambda(n)$ from $\mathcal{L}(1)$ to $\mathcal{L}(0)$ such that

### v2

# The calculus of Duistermaat’s triple index

@@ 1. Introduction
An important problem in the spectral theory of metric graphs is to quantify changes in the Laplacian eigenvalues when a finite number of vertex conditions are altered. [^1] For a survey of results when the vertex conditions are of a particularly well-studied type, we refer the reader to [6, Sec. 3.1]. Abstractly, this amounts to comparing self-adjoint extensions of a symmetric operator with finite and equal defect numbers. It was recently observed [5] that the difference of eigenvalue counting functions can be efficiently bounded in terms of a certain topological index.
The relevant index, introduced by Duistermaat in [10], takes three Lagrangian subspaces as its input. In the metric graph setting these correspond to the two sets of vertex conditions under consideration and also the Dirichlet conditions at the affected vertices. Duistermaat’s definition is somewhat indirect, as it involves an arbitrary choice of a fourth Lagrangian plane that is transversal to the three planes of interest. The proof that the resulting index is independent of this choice relies on the Maslov index for paths of Lagrangian subspaces, see [10, Eq. (2.16)]. The Maslov index is an invariant depending on the homotopy class of the path (not just its endpoints), and hence it is inherently more complicated than an index depending on fixed Lagrangian subspaces.
@@
The above index takes values between 0 and $n$:
@@
To motivate the terminology in (3), we recall the coboundary operator $\delta$ (see [13, Chapter 3]) which maps functions of k1k-1 arguments to functions of $k$ argument by
@@
Example 1.5.
In the symplectic space $\mathbb{C}\oplus\mathbb{C}$, consider the Lagrangian planes $\mathcal{G}_{A}$, $\mathcal{G}_{B}$, and $\mathcal{G}_{C}$, where $A,B,C\in\mathbb{R}$ and $\mathcal{G}_{H}:=\{(z,Hz)\colon z\in\mathbb{C}\}$. Theorem 1.4 immediately yields
[…]
These results (with distinct $A$, $B$ and $C$) are illustrated in Figure 1.
Figure 1. Possible transversal configurations of three Lagrangian planes (illustrated in $\mathbb{R}^{2}$) and the corresponding values of the Duistermaat index, see Example 1.5. The directed dashed arcs indicate the non-decreasing paths from $A$ to $B$, realizing the minimum in (1.15).
The plan for the rest of the note is as follows: in Section 2 we show that an index satisfying the conditions in Theorem 1.1 exists and in Section 3 we establish some further properties of this index. In Section 4 we interpret the Duistermaat index as the minimal value of the Maslov index [^2] Our definition of the Maslov index, which is reviewed in Section 4, follows the normalization convention of [24]; see [5, Sec. A.4] for a comparison with other conventions in the literature. over non-decreasing paths, namely
@@
As a corollary, see Theorem 4.8, we obtain a short proof of the Zhou–Wu–Zhu identity [24, Thm. 1.1], which expresses the so-called Hörmander index as a difference of Duistermaat indices. Finally, in Section 5, we prove Theorem 1.4, which we use to deduce (1.1) and its extensions from equation (1.2). We also give an elementary derivation of (1.1) using the Haynsworth inertia formula.
@@ Acknowledgements
This material is based upon work (G.B.) supported by a grant from the Institute for Advanced Study (IAS), School of Mathematics. G.B. also acknowledges support from the NSF Grants DMS-2247473, DMS–2510345 and from the Association of Former Students of Texas A&M University (through the Faculty Development Leave program that supported the leave to IAS). G.C. acknowledges the support of NSERC grant RGPIN-2017-04259. Y.L. was supported by the NSF grant DMS-2106157, and would like to thank the Courant Institute of Mathematical Sciences at NYU and especially Prof. Lai-Sang Young for their hospitality. S.S. was supported in part by the NSF Grants DMS-2418900 and DMS–2510344, Simons Foundation grant MP–TSM–00002897, grant no. $2024154$ from the U.S.-Israel Binational Science Foundation Jerusalem, Israel, and by the Office of the Vice President for Research & Economic Development (OVPRED) at Auburn University through the Research Support Program grant.
@@ 2. Existence of the Duistermaat index
[…]
@@
$R^{\epsilon}:=Y(X+\epsilon Y)^{-1},\qquad\epsilon\in\mathbb{R}\backslash E(\mathcal{M}),$ (2.5)
@@
[…]
@@
We follow the proof of a similar identity in [20, Lem. 8.3] (attributed therein to Brian Feldstein). We first compute the form $\Omega$ defined in (2.1) as a Hermitian form on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}\oplus\mathbb{C}^{n}$, where each $\mathbb{C}^{n}$ parametrizes the corresponding Lagrangian plane $\mathcal{L}_{j}$ via a frame $(X_{j},Y_{j})$. If, for instance,
@@
$=\big\{v\in\mathbb{C}^{n}\colon 0=\left<u,(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k})v\right>_{\mathbb{C}^{n}}\text{ for all }u\in\mathbb{C}^{n}\big\}$
@@
Finally, to estimate the range of the possible values of the index, we use a symplectic transformation $\mathsf{S}$ such that $\mathsf{S}\mathcal{L}_{3}=\mathcal{G}_{\infty}$ to reduce the question to bounding the range of $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{G}_{\infty})$. That the latter is between 0 and $n$ follows from a version of equation (2.7):
@@
Equation (2.21) was already established in [5, Eq. (3.32)], but we repeat the proof here for completeness. We use equation (2.7) with $R^{\epsilon}_{3}=\epsilon^{-1}I$, and the desired result follows once we establish that $n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)=0=n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)$. In other words, we claim that for any Lagrangian frame $(X,Y)^{T}$, the Hermitian operator $\epsilon^{-1}I-Y(X+\epsilon Y)^{-1}$ is non-negative definite for small enough $\epsilon>0$. It is equivalent to consider
@@ 3.1. General identities
$0$ $=(\delta\!\operatorname{\iota})(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ (3.9)
@@ 3.2. Self-adjoint linear relations
For a given splitting of the symplectic space (in our case, $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$), every Lagrangian plane can be viewed as a self-adjoint linear relation; see [22, Sec. 14.1] or [8, Sec. 4.2]. In this context, the difference of two Lagrangian planes, $\mathcal{L}$ and $\mathcal{M}$, is the Lagrangian plane
$\mathcal{L}-\mathcal{M}:=\left\{\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}-v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}\end{pmatrix}\in\mathcal{L},\ \begin{pmatrix}u\\ v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathcal{M}\right\}$, (3.11)
@@
$\operatorname{dom}\mathcal{L}$ $:=\left\{u\in\mathbb{C}^{n}:\begin{pmatrix}u\\ v\end{pmatrix}\in\mathcal{L}\text{ for some }v\in\mathbb{C}^{n}\right\},$ (3.13)
$\operatorname{mul}\mathcal{L}$ $:=\left\{v\in\mathbb{C}^{n}:\begin{pmatrix}0\\ v\end{pmatrix}\in\mathcal{L}\right\}$, (3.14)
@@
$\mathcal{L}=\big\{(x,Lx+y)\colon x\in\operatorname{dom}\mathcal{L},\ y\in\operatorname{mul}\mathcal{L}\big\}$ (3.15)
for some Hermitian matrix $L\colon\operatorname{dom}\mathcal{L}\to\operatorname{dom}\mathcal{L}$, see, e.g. [3, Thm. 1.5.1] or [22, Prop. 14.2].
@@
$\mathcal{L}-\mathcal{G}_{A}$ $=\big\{(x,Lx-Ax+y)\colon x\in\operatorname{dom}\mathcal{L},\ y\in\operatorname{mul}\mathcal{L}\big\}$ (3.22)
$=\big\{(x,(L-PA)x+y^{\prime})\colon x\in\operatorname{dom}\mathcal{L},\ y^{\prime}\in\operatorname{mul}\mathcal{L}\big\}.$
@@ 4.1. The Kashiwara index
Another commonly used index of a triple of Lagrangian planes arose independently in the work of Hörmander, Kashiwara and Wall; see, for instance, [23], [15, Eq. (3.3.15)], [17, Def. 1.5.1], [9, Sec. 8], [12, Sec 4.2]. The Hörmander–Kashiwara–Wall index $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ is defined [^4] Only the latter reference considered complex symplectic space; the signs in our definition of the form $\Omega$ are different. as the signature of the form $\Omega$ from Theorem 2.1,
@@
Proposition 4.2 is well-known when (4.7) is replaced by the condition $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=s(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$, see, e.g. [20, Thm. 8.4(d)] or [1, Prop. 4.4]. But the indices $s$ and $\operatorname{\iota}$ are related by an invertible linear transformation involving the invariants appearing in (4.5). ∎
@@
While usually established before the more complicated case of a triple, see [20, Exercise 8.1], this fact also follows from Proposition 4.2 by applying it to the triples $(\mathcal{L}_{1},\mathcal{L}_{1},\mathcal{L}_{2})$ and $(\mathcal{M}_{1},\mathcal{M}_{1},\mathcal{M}_{2})$ and using (3.1). ∎
@@ 4.2. The Maslov index
Suppose $\mathcal{L}\colon[0,1]\to\Lambda(n)$ is a differentiable path of Lagrangian planes, given by a frame $\big(X(t),Y(t)\big)$. For $t_{0}\in[0,1]$ and $v\in\mathcal{L}(t_{0})$ we define the [^5] The equivalence of (4.9) to other common formulas in the literature is proved in [5, Thm. 2.1]. crossing form $\mathfrak{q}$ on $\mathcal{L}(t_{0})$ by
@@
To calculate this Maslov index we define the two-parameter family $\mathcal{L}(s,t)=\mathcal{G}_{tQ-sI}$. It is clear that $\mathcal{L}(t)=\mathcal{L}(0,t)$ is homotopic, with fixed endpoints, to the concatenation of $\mathcal{L}(s,0)|_{s\in[0,s_{*}]}$, $\mathcal{L}(s_{*},t)|_{t\in[0,1]}$ and $\mathcal{L}(s_{*}-s,1)|_{s\in[0,s_{*}]}$, so the index of $\mathcal{L}(t)$ is the sum of these three indices. We choose the parameter $s_{*}>0$ to be small enough to guarantee that neither $M$ nor $M-Q_{\operatorname{dom}\mathcal{M}}$ has any spectrum in the interval $s_{*},0)$.
@@
The path $\mathcal{L}(s,0)|_{[0,s_{*}]}$ may still have a conjugate point at $s=0$. However, the $s$-crossing form, computed via (4.9), is $\left<x,x\right>_{\mathbb{C}^{n}}$, which is negative definite and hence generates no contribution to (4.10) at the initial point of a path. We conclude that $\operatorname{Mas}_{[0,s_{*}]}\!\big(\mathcal{L}(s,0),\mathcal{M}\big)=0$ and, by similar reasoning, $\operatorname{Mas}_{[0,s_{*}]}\!\big(\mathcal{L}(s_{*}-s,1),\mathcal{M}\big)=0$.
We now evaluate the remaining term, $\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)$. The $t$-crossing form evaluates to $\left<x,Qx\right>_{\mathbb{C}^{n}}$, which we claim to be strictly positive on $\mathcal{L}(s_{*},t)\cap\mathcal{M}$ whenever this intersection is non-trivial. Indeed, the form is obviously non-negative and if $Qx=0$ for some nonzero $x\in\operatorname{dom}\mathcal{L}$, then (4.16) implies $Mx=-s_{*}x$, which is excluded by our condition that $M$ has no spectrum in $s_{*},0)$.
@@
As a corollary of Lemma 4.6 we obtain the Zhou–Wu–Zhu identity [24, Thm. 1.1], which has numerous applications in the analysis of differential equations [2, 16, 11, 5, 18].
@@
Use Lemma 4.6 to choose a path $\widetilde{\mathcal{L}}\colon[1,2]\to\Lambda(n)$ from $\mathcal{L}(1)$ to $\mathcal{L}(0)$ such that